mary is y year old.her father is 4 times older then mary.her mother is 7 years younger than her father.if their three ages add up to 101 years,find the value of y. find the ages of mary's parents

Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

mary is y year old.her father is 4 times older then mary.her mother is 7 years younger than her father.if their three ages add up to 101 years,find the value of y. find the ages of mary's parents
-------
Mary: y years
Father: 4y years
Mother: 4y-7 years
----
Equation:
y + 4y + 4y-7 = 101
9y = 94
y = 10.44 years (Mary's age)
4y = 41.78 years (Father's age)
4y-7 = 34.78 years (Mother's age)
=====================================
Cheers,
Stan H.
==================

Answer by checkley79(3341)

MARY=Y
FATHER=4Y
MOTHER=4Y-7
Y+4Y+4Y-7=101
9Y=101+7
9Y=108
Y=108/9
Y=12 AGE OF THE DAUGHTER.
4*12=48 AGE OF THE FATHER.
48-7=41 AGE OF THE MOTHER.
PROOF:
12+48+41=101
101=101